OpenStudy (anonymous):

SUPER EASY, QUICK Navigate questions by number 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 Submit Test Question #5 / 30 The yearbook club had a meeting. The meeting had 30 people, which is five-sixths of the club. How many people are in the club?

OpenStudy (anonymous):

I just blanked out, what do I do? Divide 6?

OpenStudy (mertsj):

5/6 of the club = the people at the meeting

OpenStudy (anonymous):

5/6=30 5=180

OpenStudy (anonymous):

divide 30 by 5 and you get 6 30 + 6 =

OpenStudy (anonymous):

It's 36 THanks guys I"m so dumb

OpenStudy (anonymous):

Your welcome :)